OpenStudy (xochital):

Find the value of 10 times the square root of 2 over the square root of 5..

OpenStudy (anonymous):

lol calculate that

OpenStudy (xochital):

could you explain how to do that?

OpenStudy (anonymous):

\[\frac{ 10 \sqrt2 }{ \sqrt5 }=\frac{ 5*2*\sqrt2 }{ \sqrt5 }=\frac{ \sqrt5*\sqrt5*\sqrt2 }{ \sqrt5 }=\sqrt5*\sqrt2=\sqrt{10}\] there you go babe ;)